OpenStudy (anonymous):

suppose a and b are positive integers. Prove if ab is odd, then both a and b are odd.

OpenStudy (anonymous):

If all you're doing is proving it that a and b are odd, then substitute in odd numbers for a and b. 3*3= 9 5*7=35 9*9=81

OpenStudy (anonymous):

I have to do it theoretically not with numbers.